# Differences in Typological Alignment in Language Models’ Treatment of Differential Argument Marking

### v1

@@ 1 Introduction
Recent years have seen growing interest in whether language models (LMs) exhibit typological tendencies that resemble cross-linguistic regularities observed in human languages (Kallini et al., 2024; Xu et al., 2025). A prominent line of work addressing this question adopts the synthetic corpus paradigm, which allows researchers to systematically compare grammatical systems, linguistic features, and learning conditions using artificial corpora or modified natural corpora (Kajikawa et al., 2024; Patil et al., 2024; Leong and Linzen, 2024; Yao et al., 2025). Using this paradigm, previous studies have primarily focused on structural properties of grammar, such as word order and dependency configurations, and have shown that LMs can develop non-trivial generalizations that exhibit partial alignment with well-known typological universals like word order universals (Kuribayashi et al., 2024; Xu et al., 2025; El-Naggar et al., 2025). However, it remains unclear whether the typological tendencies observed under controlled training extend beyond purely structural phenomena to systems where grammatical well-formedness depends on semantic factors. Differential argument marking (DAM) provides a natural and revealing test case.
Figure 1: Overview of the DAM rule injection and the overall results for markedness and argument preference averaged from all DAM rule mastery accuracies.
@@
In this paper, we ask whether LMs trained under supervised learning conditions exhibit typological biases in DAM, and if so, whether such biases align with those observed in human languages. To address this question, we adopt the synthetic learning paradigm. We train GPT-2-small (Radford et al., 2019) models from scratch on 18 corpora where we perturb the SVO sentences in natural English text to implement different DAM systems. We test how well models master the DAM rules using minimal pairs and evaluate model performance across typological conditions. Our results reveal a dissociation between different types of typological asymmetries. Models consistently reproduce typological preferences in markedness, favoring systems in which marked forms align with less usual configurations, consistent with the typological patterns. Meanwhile, models diverge from human languages in their argument preferences, not showing object preference. Figure 1 provides an overview of the DAM rule injection and summarizes the overall results.
We argue that this selective alignment provides evidence about the sources of different typological asymmetries. Our findings suggest that natural markedness can emerge from distributional regularities and formal learnability, consistent with accounts that derive markedness from structural prominence (Aissen, 2003). In contrast, the failure to reproduce object preference indicates that this asymmetry may depend on discourse structure, thematic prominence, and communicative pressures not captured by standard next-token training (Iemmolo, 2010; Tal et al., 2022). Taken together, these results suggest that distinct typological tendencies reflect different underlying pressures, some of which are accessible to LMs while others are not.
@@ 2.1 Synthetic Corpus Paradigms and Typological Tendencies in LMs
Synthetic corpus paradigms construct artificial languages or modify natural corpora to enable controlled comparison of grammatical systems, linguistic features, and learning conditions. Prior work has used them to compare communicative efficiency (Kajikawa et al., 2024), test model generalization of target phenomenon from indirect distributional evidence (Patil et al., 2024; Leong and Linzen, 2024; Yao et al., 2025), and examine how semantic or pragmatic cues shape learning preferences (Misra and Kim, 2024).
@@ 2.2 Differential Argument Marking
A second typological asymmetry in DAM concerns argument preference, that is, whether the subject or object more frequently receives overt marking. Cross-linguistically, object-targeting systems overwhelmingly predominate, while systems that differentially mark subjects are much rarer (Seržant2018). Despite their relative rarity, subject marking systems remain theoretically important: recent comparative work proposes that differential subject marking can be analyzed as functionally parallel to object marking, even though they are less frequently seen (Just, 2024).
Argument preference is typically explained by communicative and discourse pressures. Iemmolo (2010) argues that overt object marking originates as a discourse-pragmatic strategy: objects are canonically less likely to function as primary topics, so overt marking arises when they assume atypical discourse roles such as topicality, signaling a departure from the default subject–topic alignment. Over time, this discourse-conditioned marking may grammaticalize into stable, object-centered systems, whereas subjects commonly aligned with topicality less likely require additional marking.
@@ 4.1 Corpus Construction
We parse the corpus using spaCy (Honnibal and Montani, 2017), augmented with Benepar constituency parsing (Kitaev and Klein, 2018), to extract SVO clauses and automatically annotate argument-level semantic properties. Semantic triggers including animacy, definiteness, and pronominality are determined using a BERT-based binary classifier trained on  2k NP instances from the validation split labeled for their semantic properties. Based on these annotations, we inject DAM by inserting special marker tokens at the right edge of the licensed argument whenever the corresponding rule-specific condition is met. Detailed corpus statistics, preprocessing steps, annotation methods and human inspection are provided in Appendix A.
@@ Results.
Figure 2 shows learning curves for rule mastery accuracy across training, while Figure 3 reports the best accuracy for each DAM condition. Overall, LMs exhibit several consistent patterns in DAM rule mastery across different dimensions.
@@
Second, within local conditions, natural rules are consistently learned better than their inverse counterparts (averaged best accuracy: natural $\approx$ 0.85 vs. inverse $\approx$ 0.68). This advantage is visible in Figure 3, where natural variants outperform inverse variants across semantic triggers in both Local-A and Local-P. The same pattern holds throughout training in Figures 2(a) and 2(b), where solid lines (natural) consistently remain above dashed lines (inverse) of the same color. As for argument preference, we do not observe a clear advantage of Local-P over Local-A in rule mastery accuracy in Figure 3 (averaged best accuracy: subject-targeting $\approx$ 0.74 vs. object-targeting $\approx$ 0.79). We will discuss these asymmetries in detail in Section 5.
Third, we observe that the performance gap between natural and inverse variants is smaller for Local-P rules than for Local-A rules. One possible explanation is that the model encodes semantic properties of subjects more robustly than those of objects. Marginal distributions reported in Table 5 in Appendix A.3 show that prominent arguments occur more frequently in subject position than low-prominence arguments occur in object position. Semantic probing analyses reported in Appendix C.1 further indicate that the model encodes subject-related semantic features more strongly and stably.
@@
To examine whether rule mastery is driven primarily by input frequency, we analyze the relationship between each rule’s best accuracy and its perturbed ratio in the training corpus (SVO% in Table 3). Across all rules, perturbed ratio and best accuracy exhibit a significant negative correlation (r=0.56r=-0.56, $p=0.0166$).
However, this negative correlation is largely driven by differences between natural and inverse rules. Within the subset of high-performing Local–Natural rules, where overall accuracy is the highest, the correlation between perturbed ratio and accuracy largely disappears (r=0.17r=-0.17, $p=0.7536$). These results indicate that a lower perturbed ratio does not systematically lead to better rule mastery, and that exposure alone is insufficient to explain differences in learnability across DAM rules.
@@ 5 Conclusion and Discussion
We discussed two DAM typological tendencies in LMs: (i) markedness direction, i.e., whether marking occurs on less usual arguments, and (ii) argument preference, i.e. whether marking targets the subject or object. For markedness direction, we find a clear and consistent advantage for natural systems over inverse systems, aligning with DAM typological tendencies in human language. For argument preference, we do not find object-targeting rules outperforming subject-targeting rules, which does not align with the typological tendency.
@@
It is important to note that while markedness and argument preference may be influenced by different mechanisms, these mechanisms are not mutually exclusive, nor are they the sole cause for an LM typological tendency. Prior work shows that markedness can emerge through communicative pressures even from initially random case-marking systems (Fedzechkina et al., 2012, 2017; Smith and Culbertson, 2025; Lian et al., 2025). Yet our models, trained solely under supervised next-token prediction, still display typologically consistent markedness patterns. This suggests that markedness cannot be reduced to communicative efficiency or predictability alone, but instead reflects the joint influence of learnability and communicative pressures.
@@ 6 Limitations
Our experiments are conducted on relatively small models and datasets, leaving open whether the observed patterns persist at larger scales. Also, we report results from a single random seed; evaluating across multiple seeds would provide a more reliable estimate of performance. Beyond this, we also see three methodological aspects that are worth improving.
@@
Second, we perturb only clauses with a single finite verb, one subject, and one nominal direct object, leaving out ditransitives, passives, clausal complements, and other construction types. However, in natural languages, case marking often interacts with a broader range of argument-structure alternations. Future work could extend the paradigm to a richer and more structurally diverse set of clause types.
Third, our corpus construction strategy prioritizes preserving the overall distribution of argument structures and semantic properties present in the original data. This design choice allows us to study DAM learning under relatively naturalistic input distributions. However, it also means that the triggering frequency of different DAM rules is not experimentally controlled. As a result, while we analyze the relationship between input frequency and rule mastery, the current design cannot fully disentangle frequency effects from structural learning constraints. Future work could address this limitation by explicitly controlling the frequency of different SVO configurations, enabling a more precise separation of frequency-driven effects from inherent learnability differences across rules.
@@ A.2 Parsing and SVO Extraction
Each sentence is parsed using spaCy (Honnibal and Montani, 2017), augmented with Benepar constituency parsing (Kitaev and Klein, 2018). For each sentence, we identify verbal predicates and extract clause-local predicate–argument frames by considering only dependents that are directly licensed by the predicate head. Each frame consists of one predicate together with at most one subject and one object-like argument per clause. Instances with multiple object candidates are excluded.
@@
In addition to bare nominal objects, we treat certain prepositional constructions as pseudo-objects when they function as predicate-selected complements. Such constructions are grouped into a single argumental unit spanning the preposition and its nominal complement. This design reflects a typological and constructional property of English, in which patient-like roles are frequently encoded via prepositional marking (e.g., wait for the bus, listen to the story) rather than bare object positions. Cross-linguistically, such prepositional realizations often correspond to bare objects in languages with richer case morphology, motivating their inclusion in our case-marking framework (Blake, 2001).
@@ A.4 Rule Injection and Dataset Construction
In addition to the 18 experimental conditions, we include two control settings for comparison: (i) an unperturbed baseline trained on the original corpus, and (ii) a fully perturbed condition that inserts both agent and patient markers on every eligible S–V–O frame irrespective of licensing. Each marker is inserted at the right edge of the recorded NP span, [^3] In some languages, case markers are realized on the head of a phrase rather than at its edge. We do not model such head-marking configurations in the current setup. preserving all original tokens and punctuation.
@@ A.5 Human Inspection
To assess the reliability of the perturbation pipeline, we conducted a targeted human inspection on a small set of representative DAM rules. Specifically, we inspected L-P-Def, L-A-Pro-inv, and G-Ani, which together cover different dimensions of the design space. For each rule, we manually examined 50 affected and 50 unaffected sentences. In addition, we inspected 100 sentences sampled from the shared Invalid pool. The results are summarized in Table 6.
@@ Appendix B Model Training Details
Epochs $\sim$8.0–8.2444Models are trained for a fixed number of optimization steps. The effective number of epochs varies slightly across DAM rules due to differences in marker density.
@@ C.1 Semantic Probing
We conduct a semantic probing experiment to test whether argument-level semantic properties are linearly recoverable from model representations using standard probing techniques (Veldhoen et al., 2016; Hewitt and Liang, 2019; Belinkov, 2022). We focus on three binary features underlying our DAM rules: animacy, definiteness, and pronominality.

### v2

@@ 1 Introduction
Recent years have seen growing interest in whether language models (LMs) exhibit typological tendencies that resemble cross-linguistic regularities observed in human languages (Kallini et al., 2024; Xu et al., 2025). A prominent line of work addressing this question adopts the synthetic corpus paradigm, which allows researchers to systematically compare grammatical systems, linguistic features, and learning conditions using artificial corpora or modified natural corpora (Kajikawa et al., 2024; Patil et al., 2024; Leong and Linzen, 2026; Yao et al., 2025). Using this paradigm, previous studies have primarily focused on structural properties of grammar, such as word order and dependency configurations, and have shown that LMs can develop non-trivial generalizations that exhibit partial alignment with well-known typological universals like word order universals (Kuribayashi et al., 2024; Xu et al., 2025; El-Naggar et al., 2025). However, it remains unclear whether the typological tendencies observed under controlled training extend beyond purely structural phenomena to systems where grammatical well-formedness depends on semantic factors. Differential argument marking (DAM) provides a natural and revealing test case.
Figure 1: Overview of the DAM rule injection and the overall results for markedness and argument preference averaged over all DAM rule mastery accuracies.
@@
In this paper, we ask whether LMs trained with a standard next-token prediction objective exhibit typological biases in DAM, and whether such biases align with those observed in human languages. We examine this question in a controlled synthetic-learning setup using small autoregressive models. Specifically, we train GPT-2-small (Radford et al., 2019) models from scratch on 18 corpora, where we perturb SVO sentences in natural English text to implement different DAM systems. We test how well models master the DAM rules using minimal pairs and evaluate model performance across typological conditions. Our results reveal a dissociation between different types of typological asymmetries. Models consistently reproduce typological preferences in markedness, favoring systems in which marked forms align with less usual configurations, consistent with the typological patterns. Meanwhile, models diverge from human languages in their argument preferences, not showing a strong object preference. Figure 1 provides an overview of the DAM rule injection and summarizes the overall results.
We argue that this selective alignment provides evidence about the sources of different typological asymmetries. Our findings suggest that natural markedness can emerge from distributional regularities and formal learnability, consistent with accounts that derive markedness from structural prominence (Aissen, 2003). In contrast, the absence of a strong object preference indicates that this asymmetry may depend on discourse structure, thematic prominence, and communicative pressures not captured by standard next-token training (Iemmolo, 2010; Tal et al., 2022). Taken together, these results suggest that distinct typological tendencies reflect different underlying pressures, some of which are accessible to LMs while others are not.
@@ 2.1 Synthetic Corpus Paradigms and Typological Tendencies in LMs
Synthetic corpus paradigms construct artificial languages or modify natural corpora to enable controlled comparison of grammatical systems, linguistic features, and learning conditions. Prior work has used them to compare communicative efficiency (Kajikawa et al., 2024), test model generalization of target phenomenon from indirect distributional evidence (Patil et al., 2024; Leong and Linzen, 2026; Yao et al., 2025), and examine how semantic or pragmatic cues shape learning preferences (Misra and Kim, 2024).
@@ 2.2 Differential Argument Marking
A second typological asymmetry in DAM concerns argument preference, that is, whether the subject or object more frequently receives overt marking. Cross-linguistically, object-targeting systems overwhelmingly predominate, while systems that differentially mark subjects are much rarer (Schmidtke-Bode and Levshina, 2018; Seržant2018). Despite their relative rarity, subject marking systems remain theoretically important: recent comparative work proposes that differential subject marking can be analyzed as functionally parallel to object marking, even though they are less frequently seen (Just, 2024).
Argument preference is typically explained by communicative and discourse pressures. Iemmolo (2010) argues that overt object marking originates as a discourse-pragmatic strategy: objects are canonically less likely to function as primary topics, so overt marking arises when they assume atypical discourse roles such as topicality, signaling a departure from the default subject–topic alignment. Over time, this discourse-conditioned marking may grammaticalize into stable, object-centered systems, whereas subjects commonly aligned with topicality are less likely to require additional marking.
@@ 4.1 Corpus Construction
We parse the corpus using spaCy (Honnibal et al., 2020), augmented with Benepar constituency parsing (Kitaev and Klein, 2018), to extract SVO clauses and automatically annotate argument-level semantic properties. Semantic triggers, including animacy, definiteness, and pronominality, are determined using a BERT-based binary classifier trained on $\sim$2k NP instances from the validation split labeled for their semantic properties. Based on these annotations, we inject DAM by inserting special marker tokens at the right edge of the licensed argument whenever the corresponding rule-specific condition is met. We realize these markers as independent tokens to avoid introducing additional learning noise from BPE segmentation. Detailed corpus statistics, preprocessing steps, annotation methods and human inspection are provided in Appendix A.
@@ Results.
Figure 2 shows learning curves for rule mastery accuracy across training, while Figure 3 reports the best accuracy for each DAM condition. Overall, our models exhibit several consistent patterns in DAM rule mastery across different dimensions.
@@
Second, within local conditions, natural rules are consistently learned better than their inverse counterparts (averaged best accuracy: natural $\approx$ 0.85 vs. inverse $\approx$ 0.68). This advantage is visible in Figure 3, where natural variants outperform inverse variants across semantic triggers in both Local-A and Local-P. The same pattern holds throughout training in Figures 2(a) and 2(b), where solid lines (natural) consistently remain above dashed lines (inverse) of the same color.
By contrast, the overall effect of argument target is much smaller. Although Local-P rules are slightly more accurate on average than Local-A rules (object-targeting $\approx$ 0.79 vs. subject-targeting $\approx$ 0.74), this difference is substantially smaller than the natural–inverse contrast. This pattern is further supported by a local-only ANOVA over the 12 local conditions, with direction, argument target, and semantic trigger as crossed factors. Markedness direction accounts for the largest share of variance in rule mastery accuracy ($\eta^{2}=0.566$), whereas the main effects of semantic trigger ($\eta^{2}=0.111$) and argument target ($\eta^{2}=0.050$) are comparatively small.
Third, we observe that the performance gap between natural and inverse variants is smaller for Local-P rules than for Local-A rules. In the local-only ANOVA, the direction $\times$ target interaction also explains more variance than the target main effect ($\eta^{2}=0.106$ vs. $0.050$). One possible explanation is that the model encodes semantic properties of subjects more robustly than those of objects. Marginal distributions reported in Table 5 in Appendix A.3 show that prominent arguments occur more frequently in subject position than low-prominence arguments occur in object position. Semantic probing analyses reported in Appendix C.1 further indicate that the model encodes subject-related semantic features more strongly and stably.
@@
Figure 4: Scatter plot of the perturbation ratio in the training corpus (ALL%) versus best rule mastery accuracy across DAM rules. The dashed line shows the linear regression fit.
@@
To examine whether differences in DAM rule mastery can be explained primarily by input frequency, we analyze model performance as a function of perturbed ratio in the training corpus (ALL% in Table 3), as shown in Figure 4. Across all rules, perturbed ratio and rule mastery exhibit a moderate negative correlation (r=0.56r=-0.56, $p=0.0166$).
However, frequency alone does not fully account for the observed asymmetries in model performance. While natural rules (blue) tend to cluster around higher performance and lower perturbation ratios and inverse rules (red) around lower performance and higher perturbation ratios, substantial variation remains within each group. Within the subset of high-performing Local–Natural rules, the correlation between perturbed ratio and accuracy largely disappears (r=0.17r=-0.17, $p=0.7536$), suggesting that lower perturbation ratios do not systematically lead to better rule mastery.
At the same time, markedness direction and frequency are inherently linked in naturalistic distributions: overt marking in natural languages is typically associated with less frequent or less expected argument configurations. As a result, manipulating the proportion of marked and unmarked instances also changes the typological properties of the system itself. We therefore treat frequency as a relevant factor in rule mastery, but not as a sufficient explanation for the observed pattern.
@@ 5 Conclusion and Discussion
We discuss two DAM typological tendencies in GPT-2-small models: (i) markedness direction, i.e., whether marking occurs on less usual arguments, and (ii) argument preference, i.e. whether marking targets the subject or object. For markedness direction, we find a clear and consistent advantage for natural systems over inverse systems, aligning with DAM typological tendencies in human language. For argument preference, object-targeting rules show only a weak overall advantage over subject-targeting rules, which does not align with the typological tendency.
@@
Because our corpora are constructed by injecting DAM rules into English text, distributional and morphosyntactic cues from the underlying corpus may also shape model behavior. For markedness direction, English already contains broader markedness patterns, such as unmarked singulars versus overtly marked plurals. The advantage of natural DAM rules may therefore reflect their consistency with this broader markedness direction, rather than a purely DAM-specific preference. For argument preference, English retains subject–verb agreement as a head-marking cue, which may support learning subject-targeting DAM rules. These residual cues could in principle affect model preferences if they were controlled or removed. We treat this as an important limitation, reflecting a broader constraint of the current corpus perturbation methodology.
It is important to note that while markedness and argument preference may be influenced by different mechanisms, these mechanisms are not mutually exclusive, nor are they the sole cause for an LM typological tendency. Prior work shows that markedness can emerge through communicative pressures even from initially random case-marking systems (Fedzechkina et al., 2012, 2017; Smith and Culbertson, 2025; Lian et al., 2025). Yet our models, trained solely with a standard next-token prediction objective, still display typologically consistent markedness patterns. This suggests that markedness cannot be reduced to communicative efficiency or predictability alone, but instead reflects the joint influence of learnability and communicative pressures.
@@ 6 Limitations
Our experiments are conducted with a single small autoregressive architecture, GPT-2-small, and on a fixed data scale, leaving open whether the observed patterns persist across model families, larger model sizes, or different training regimes. We also report results from a single random seed; evaluating across multiple seeds would provide a more reliable estimate of performance. Beyond this, we also see three methodological aspects that are worth improving.
@@
Second, we perturb only clauses with a single finite verb, one subject, and one nominal direct object, leaving out ditransitives, passives, raising and control structures, and other construction types. However, in natural languages, case marking often interacts with a broader range of argument-structure alternations. Future work could extend the paradigm to a richer and more structurally diverse set of clause types.
Third, our implementation captures only a simplified version of the DAM design space. We operationalize animacy, definiteness, and pronominality as separate binary triggers, but these dimensions are not fully independent in natural languages. For example, work on Spanish DOM argues that animacy is the dominant trigger of object marking, while apparent specificity effects are closely tied to topicality and information structure (Leonetti, 2004). On the morphological side, our implementation uses a single independent marker at the right edge of each marked NP, abstracting away from multiple coding devices, morphologically integrated case marking, and marking on nominal heads (Seržant2018). Our setup also treats triggers as sentence-internal properties, leaving aside discourse-sensitive DAM systems where marking depends on information structure, topic continuity, or context-dependent shifts in prominence or animacy.
Finally, our corpus construction strategy prioritizes preserving the overall distribution of argument structures and semantic properties present in the original data. This design choice allows us to study DAM learning under relatively naturalistic input distributions. However, it also means that the triggering frequency of different DAM rules is not experimentally controlled. As a result, while we analyze the relationship between input frequency and rule mastery, the current design cannot fully disentangle frequency effects from structural learning constraints, as discussed in Section 4.3. Future work could address this limitation by explicitly controlling the frequency of different SVO configurations, enabling a more precise separation of frequency-driven effects from inherent learnability differences across rules.
@@ A.2 Parsing and SVO Extraction
Each sentence is parsed using spaCy (Honnibal et al., 2020), augmented with Benepar constituency parsing (Kitaev and Klein, 2018). For each sentence, we identify verbal predicates and extract clause-local predicate–argument frames by considering only dependents that are directly licensed by the predicate head. Each frame consists of one predicate together with at most one subject and one object-like argument per clause. Instances with multiple object candidates are excluded.
@@
In addition to bare nominal objects, we treat certain prepositional constructions as pseudo-objects when they function as predicate-selected complements. Operationally, we identify these cases as prepositional dependents attached to the verbal predicate whose complement forms a patient/theme-like argument. Such constructions are grouped into a single argumental unit spanning the preposition and its nominal complement. This design reflects a typological and constructional property of English, in which patient-like roles are frequently encoded via prepositional marking (e.g., wait for the bus, listen to the story) rather than bare object positions. Cross-linguistically, such prepositional realizations often correspond to bare objects in languages with richer case morphology, motivating their inclusion in our case-marking framework (Blake, 2001).
@@ A.4 Rule Injection and Dataset Construction
In addition to the 18 experimental conditions, we include two control settings for comparison: (i) an unperturbed baseline trained on the original corpus, and (ii) a fully perturbed condition that inserts both agent and patient markers on every eligible S–V–O frame irrespective of licensing. Each marker is inserted at the right edge of the recorded NP span, preserving all original tokens and punctuation.
@@ A.5 Human Inspection
To assess the reliability of the perturbation pipeline, we conduct a targeted human inspection on a small set of representative DAM rules. Specifically, we inspect L-P-Def, L-A-Pro-inv, and G-Ani, which together cover different dimensions of the design space. For each rule, we manually examine 50 affected and 50 unaffected sentences. In addition, we inspect 100 sentences sampled from the shared Invalid pool. The results are summarized in Table 6.
@@ Appendix B Model Training Details
Epochs $\sim$8.0–8.2333Models are trained for a fixed number of optimization steps. The effective number of epochs varies slightly across DAM rules due to differences in marker density.
@@ C.1 Semantic Probing
We conduct a semantic probing experiment to test whether argument-level semantic properties are linearly recoverable from model representations using standard probing techniques (Hupkes et al., 2018; Hewitt and Liang, 2019; Belinkov, 2022). We focus on three binary features underlying our DAM rules: animacy, definiteness, and pronominality.